OpenStudy (anonymous):

solve using the multiplication principle -6x>1/13

jimthompson5910 (jim_thompson5910):

\[\Large -6x > \frac{1}{13}\] \[\Large \left(-6x\right)\left(-\frac{1}{6}\right) < \frac{1}{13}\left(-\frac{1}{6}\right)\] \[\Large x < -\frac{1}{13*6}\] \[\Large x < -\frac{1}{78}\]

OpenStudy (anonymous):

clear as purified drinking water.Thanks a million!

jimthompson5910 (jim_thompson5910):

np